A contribution margin income statement organizes costs by behavior

Contribution Margin Income Statement

A contribution margin income statement organizes costs by behavior (variable or fixed), rather than by function (operating, selling, or administrative). The contribution margin is the difference between sales and variable expenses .

Byron Manufacturing has one product that sells for $24.00 per unit. The company estimates fixed costs at $6,000, direct materials at $4.00 per unit, direct labor at $5.00 per unit, and variable overhead costs at $3.00 per unit.

Fill in the contribution margin income statement when 730 units are sold:

Byron Manufacturing
Contribution Margin Income Statement
Sales$
Less: Variable costs
Contribution margin$
Less: Fixed costs
Operating income$

Calculate Byron Manufacturing's per unit contribution margin : $.

The contribution margin ratio is 50% .

Calculating the Break-even Point:

The break-even point in sales dollars is $ which is a break-even point in units of  units.

Feedback

Use the sell price and the number of units sold to calculate sales. Use the variable costs and the number of units sold to calculate variable costs. Fixed costs are provided. Refer to the rollovers for help with the calculations for per-unit contribution margin, contribution margin ratio, and break-even point.

CVP Analysis using a chart:

The cost-volume-profit chart for Byron Manufacturing is shown. Use the graph to complete the sentences given below.

Byron Manufacturing reaches its break-even level of activity when it sells 500  units and generates $12,000  in revenue, because at this level of activity the firm's revenue equals  its total cost. In addition, you can determine from the chart that Byron Manufacturing's fixed costs are $6,000  and its price per unit is $24.00  and variable cost per unit is $12.00 .

If fixed costs increase, what will happen to the break-even point?

The break-even point will increase.

If the price per unit decreases, what will happen to the break-even point?

The break-even point will increase.


The CVP chart illustrates the calculations done in the preceding part. The results are the same.

CVP analysis is used to analyze the effects of changes in selling prices, costs and volume on profits. It is also used to determine target profit, the margin of safety, operating leverage, product mix and choosing among marketing strategies and others.

Suppose Byron management has a target operating income of $3,000. Assume the same costs as above and the sell price remains at $24 per unit. How many units does Byron need to sell to meet this goal?


What is Byron's margin of safety in sales and in units when Byron sells 730 units?

Margin of safety in sales $ Margin of safety in units  units

What is the degree of operating leverage when 730 units are sold? If required, round your answer to two decimal places.
Answer #1

In first question, we are required to prepare contribution margin statement.

Solution:

Preparation of contribution margin statement:

Particular Working Amount($)
Sales 730*24 17520
Less: Variable cost
a) Material 730*4= 2920
b) Labour 730*5= 3650
c) Overhead 730*3= 2190 (8,760)
Contribution Margin 8,760
Less: Fixed Cost (6,000)
Operating Income

2,760

Contribution margin per unit:

Particular Amount ($)
Sale price per unit 24
Less: Variable cost per unit
a) Material. 4
b) Labour. 5
c) Overhead. 3 (12)
Contribution Margin per unit 12

Contribution margin ratio:

= Contribution margin per unit÷ Sale price per unit

= (12÷24)×100 = 50%